# Kinetic Interacting Particle Langevin Monte Carlo

### v2

@@ B.1 Proof of Proposition 1
$\bar{\pi}(z,\bar{v})$ $\propto\exp\left(-\sum_{i=1}^{N}U(\theta,\sqrt{N}z_{i})-\frac{N}{2}\|\bar{v}\|^{2}\right).$
@@
Let us now look at $\theta$-marginal of this density, which can be written as
$\pi^{N}_{\Theta}(\theta)\propto\int_{\mathbb{R}^{Nd_{x}+d_{\theta}}}\int_{\mathbb{R}^{Nd_{x}}}e^{-\sum_{i=1}^{N}U(\theta,\sqrt{N}z_{i})-\frac{N}{2}\|\bar{v}^{z}\|^{2}}{\mathrm{d}}z_{x}{\mathrm{d}}\bar{v}.$
@@
$\pi^{N}_{\Theta}(\theta)$ $\propto{N}^{-1/2}\int_{\mathbb{R}^{Nd_{x}}}e^{-\sum_{i=1}^{N}U(\theta,x^{\prime}_{i})}{\mathrm{d}}x^{\prime},$
$\propto\left(\int_{{\mathbb{R}}^{d_{x}}}e^{-U(\theta,x^{\prime})}{\mathrm{d}}x^{\prime}\right)^{N}=\exp(N\log p_{\theta}(y)),$
since $p_{\theta}(y)=\int e^{-U(\theta,x)}{\mathrm{d}}x$ by definition where $U(\theta,x)=-\log p_{\theta}(x,y)$.
B.2 Proof of Proposition 2
Note that, by H1, $U(\theta,x)$ is jointly $\mu$-strongly convex. Let $\kappa(\theta)=-\log p_{\theta}(y)$. Using the Prekopa-Leindler inequality for strongly log-concave distributions [62, Theorem 3.8], we can see that
$\langle\theta-\theta^{\prime},\nabla\kappa(\theta)-\nabla\kappa(\theta)\rangle\geq\mu\|\theta-\theta^{\prime}\|^{2}.$
The bound then follows using an identical proof of Proposition 3 in [3].
B.3 Proof of Proposition 3
The proof follows from Lemma A.1 which shows that we can rewrite the KIPLD system as a single $d_{\theta}+Nd_{x}$-dimensional underdamped Langevin diffusion. Then with a suitable time rescaling (see Lemma 1 in [25]), we can rewrite this diffusion as a standard underdamped diffusion as shown in Lemma A.2. The result then follows from the exponential convergence of the underdamped Langevin diffusion [25, Theorem 1] with suitable modifications, e.g., see the proof of [44, Theorem 8.7]. The result then follows from the fact that $W_{2}$ between marginals are bounded by the $W_{2}$ between joint distributions.
B.4 Proof of Theorem 1
By Lemma A.2, we have the standard underdamped diffusion
@@ B.4 Proof of Theorem 1
${\mathrm{d}}\widetilde{{\mathbf{V}}}_{t}^{z}$ $=-\widetilde{\gamma}\widetilde{{\mathbf{V}}}_{t}^{z}{\mathrm{d}}t-N\nabla_{z}\bar{U}_{N}(\widetilde{{\mathbf{Z}}}_{t}){\mathrm{d}}t+\sqrt{{2{\widetilde{\gamma}}}}{\mathrm{d}}{\mathbf{B}}_{t}.$
It is apparent that the stationary measure of this scheme is
$\widetilde{\pi}(z,v)\propto e^{-N\bar{U}_{N}(z)-\frac{1}{2}\|v\|^{2}}.$
We will now look at the standard kinetic Langevin Monte Carlo discretisation of the SDE in (21) and show that this scheme coincides with KIPLMC1. Thus we can utilize the bounds in [25] directly for our scheme.
Let us write the Exponential Integrator discretisation of this scheme (abusing the notation using same letters) using the step-size $\tilde{\eta}$
$\widetilde{{\mathbf{Z}}}_{n+1}$ $=\widetilde{{\mathbf{Z}}}_{n}+\widetilde{\psi}_{1}(\tilde{\eta})\widetilde{{\mathbf{V}}}_{n}^{z}-\widetilde{\psi}_{2}(\tilde{\eta})N\nabla\bar{U}_{N}(\widetilde{{\mathbf{Z}}}_{n})+\sqrt{2\tilde{\gamma}}\tilde{\varepsilon}_{n+1},$ (22)
$\widetilde{{\mathbf{V}}}_{n+1}^{z}$ $=\widetilde{\psi}_{0}(\tilde{\eta})\widetilde{{\mathbf{V}}}_{n}^{z}-\widetilde{\psi}_{1}(\tilde{\eta})N\nabla\bar{U}_{N}(\widetilde{{\mathbf{Z}}}_{n})+\sqrt{2\tilde{\gamma}}\tilde{\varepsilon}_{n+1}^{\prime}.$
Since the function $z\mapsto N\bar{U}_{N}(z)$ is $N\mu$ strongly convex and $NL$-gradient-Lipschitz, for $\tilde{\gamma}\geq\sqrt{N\mu+NL}$ and $\tilde{\eta}\leq m/(4\tilde{\gamma}L)$, [25, Theorem 2] directly implies that
@@
$W_{2}(\nu_{n},\bar{\pi})\leq\sqrt{2}\left(1-\frac{0.75\mu\tilde{\eta}}{\tilde{\gamma}}\right)^{n}W_{2}(\nu_{0},\bar{\pi})+\frac{L\tilde{\eta}\sqrt{2d_{z}}}{\mu},$
where $\nu_{n}$ denotes the law of the discretisation (22) at time $n$ and $\tilde{\pi}$ the stationary measure given in Lemma A.3. We relate the discretisation (22) to KIPLMC1. It is straightforward to check that, after proper modifications, the discretisation (22) coincides with the KIPLMC1 with $\tilde{\eta}=\eta/\sqrt{N}$ and $\tilde{\gamma}=\sqrt{N}\gamma$. Therefore, by specializing this bound to $\theta$-dimension and substituting $\tilde{\eta}=\eta/\sqrt{N}$ and $\tilde{\gamma}=\sqrt{N}\gamma$, we obtain the convergence bound for the KIPLMC1. Combining this with the concentration result from Prop. 2, we obtain the result.
B.5 Proof of Theorem 2
Again, recall by Lemma A.2, we consider our rescaled standard underdamped diffusion, given by (21) and seek to recover from this the behaviour of KIPLMC2. Using this rescaled system we directly apply the results from [53] for our scheme.
We now write the splitting scheme for the diffusion in Lemma A.2, using step-size $\tilde{\eta}$,
$\widetilde{{\mathbf{Z}}}_{n+1}$ $=\widetilde{{\mathbf{Z}}}_{n}+\tilde{\eta}(\tilde{\delta}\widetilde{{\mathbf{V}}}^{z}_{n}+\sqrt{1-\tilde{\delta}^{2}}\varepsilon_{n+1})-\frac{\tilde{\eta}^{2}}{2}N\nabla\bar{U}_{N}(\widetilde{{\mathbf{Z}}}_{n})$ (23)
$\widetilde{{\mathbf{V}}}^{z}_{n+1}$ $=\tilde{\delta}^{2}\widetilde{{\mathbf{V}}}_{n}^{z}-\frac{\tilde{\eta}\tilde{\delta}}{2}(N\nabla\bar{U}_{N}(\widetilde{{\mathbf{Z}}}_{n})+N\nabla\bar{U}_{N}(\widetilde{{\mathbf{Z}}}_{n+1}))+\sqrt{1-\tilde{\delta}^{2}}(\tilde{\delta}\varepsilon_{n+1}+\varepsilon_{n+1}^{\prime}),$
where $\tilde{\delta}=e^{-\tilde{\gamma}\tilde{\eta}/2}$. Using the $N\,L$-gradient-Lipschitzness and $N\,\mu$ strong convexity of the function $z\mapsto N\bar{U}_{N}(z)$, and with the choice of $\tilde{\gamma}\geq 2\sqrt{NL}$ and $\tilde{\eta}\leq\frac{N\mu}{33\tilde{\gamma}^{3}}$, we apply the results of [53, Theorem 1] and [53, Proposition 11] to obtain
@@ B.5 Proof of Theorem 2
$W_{2}(\nu_{n},\bar{\pi})\leq\tilde{C}\left(\left(1-\frac{\tilde{\eta}\tilde{\mu}}{\tilde{\gamma}}\right)^{\frac{n}{2}}W_{2}(\nu_{0},\bar{\pi})+\tilde{\eta}\sqrt{2d_{z}}\frac{6\tilde{\gamma}\tilde{K}}{\tilde{\mu}}\right),$
@@
where $\nu_{n}$ is now the law of the discretisation (23) at time $n$ and $\tilde{\pi}$ the stationary measure from Lemma A.3. Further,

### v3

@@ B.2 KIPLMC2
$\widetilde{Z}_{k+1}$ $=\widetilde{Z}_{k}+\tilde{\eta}\left(\tilde{\delta}\widetilde{V}_{k}+\sqrt{1-\tilde{\delta}^{2}}\xi_{k+1}\right)-\frac{\tilde{\eta}^{2}}{2}N\nabla_{z}\bar{U}_{N}(\widetilde{Z}_{k})$
$\widetilde{V}_{k+1}$ $=\tilde{\delta}^{2}\widetilde{V}_{k}+\sqrt{1-\tilde{\delta}^{2}}(\tilde{\delta}\xi_{k+1}+\xi^{\prime}_{k+1})-\frac{\tilde{\eta}\tilde{\delta}}{2}N(\nabla_{z}\bar{U}_{N}(\widetilde{Z}_{k})+\nabla_{z}\bar{U}_{N}(\widetilde{Z}_{k+1})),$
where we recall that $\tilde{\delta}=e^{-\tilde{\gamma}\tilde{\eta}/2}$. Let us begin at looking at the discretisation coordinatewise in the scaled case,
$\widetilde{\theta}_{k+1}$ $=\widetilde{\theta}_{k}+\tilde{\eta}\left(\tilde{\delta}\widetilde{V}^{\theta}_{k}+\sqrt{1-\tilde{\delta}^{2}}\xi^{0}_{k+1}\right)-\frac{\tilde{\eta}^{2}}{2}N\nabla_{\theta}\bar{U}_{N}(\widetilde{Z}_{k}),$
$\widetilde{X}^{i}_{k+1}$ $=\widetilde{X}^{i}_{k}+\tilde{\eta}\left(\tilde{\delta}\widetilde{V}^{x_{i}}_{k}+\sqrt{1-\tilde{\delta}^{2}}\xi^{i}_{k+1}\right)-\frac{\tilde{\eta}^{2}}{2}N\nabla_{z_{i}}\bar{U}_{N}(\widetilde{Z}_{k}),$
$\widetilde{V}^{\theta}_{k+1}$ $=\tilde{\delta}^{2}\widetilde{V}^{\theta}_{k}+\sqrt{1-\tilde{\delta}^{2}}(\tilde{\delta}\xi^{0}_{k+1}+\xi_{k+1}^{0,\prime})-\frac{\tilde{\eta}\tilde{\delta}}{2}N(\nabla_{\theta}\bar{U}_{N}(\widetilde{Z}_{k})+\nabla_{\theta}\bar{U}_{N}(\widetilde{Z}_{k+1}),$
$\widetilde{V}^{x_{i}}_{k+1}$ $=\tilde{\delta}^{2}\widetilde{V}^{x_{i}}_{k}+\sqrt{1-\tilde{\delta}^{2}}(\tilde{\delta}\xi^{i}_{k+1}+\xi_{k+1}^{i,\prime})-\frac{\tilde{\eta}\tilde{\delta}}{2}N(\nabla_{z_{i}}\bar{U}_{N}(\widetilde{Z}_{k})+\nabla_{z_{i}}\bar{U}_{N}(\widetilde{Z}_{k+1})).$
We now proceed to rescaling the algorithm by observing that $\tilde{\delta}=\delta$ and recalling that $\tilde{\psi}_{0}^{\tilde{\eta}}=\psi_{0}^{\eta}$. Using this, $\tilde{\eta}=\eta/\sqrt{N}$ and $\tilde{\gamma}=\sqrt{N}\gamma$ and the identities (14), we obtain for the $\theta$-coordinates,
$\theta_{k+1}$ $=\theta_{k}+\frac{\eta}{\sqrt{N}}\left(\delta\widetilde{V}^{\theta}_{k}+\sqrt{1-\delta^{2}}\xi_{k+1}\right)-\frac{\eta^{2}}{2N}\sum_{i=1}^{N}\nabla_{\theta}U(\theta_{k},X_{k}^{i})$
$\widetilde{V}^{\theta}_{k+1}$ $=\delta^{2}\widetilde{V}^{\theta}_{k}+\sqrt{1-\delta^{2}}(\delta\xi_{k+1}+\xi^{\prime}_{k+1})-\frac{\eta\delta}{2\sqrt{N}}\sum_{i=1}^{N}\bigg(\nabla_{\theta}U(\theta_{k},X_{k}^{i})+\nabla_{\theta}U(\theta_{k+1},X_{k+1}^{i})\bigg),$
Substituting in the relationship $\widetilde{V}_{k}^{\theta}=\sqrt{N}V_{k}^{\theta}$ and $\widetilde{X}_{k}^{i}=X_{k}^{i}/\sqrt{N}$, we obtain,
$\theta_{k+1}$ $=\theta_{k}+\eta\left(\delta V^{\theta}_{k}+\sqrt{\frac{1-\delta^{2}}{N}}\xi_{k+1}\right)-\frac{\eta^{2}}{2N}\sum_{i=1}^{N}\nabla_{\theta}U(\theta_{k},X_{k}^{i})$
$V^{\theta}_{k+1}$ $=\delta^{2}V^{\theta}_{k}+\sqrt{\frac{1-\delta^{2}}{N}}(\delta\xi_{k+1}+\xi^{\prime}_{k+1})-\frac{\eta\delta}{2N}\sum_{i=1}^{N}\bigg(\nabla_{\theta}U(\theta_{k},X_{k}^{i})+\nabla_{\theta}U(\theta_{k+1},X_{k+1}^{i})\bigg).$
For the particles $i\in[N]$, the same substitutions give,
$\widetilde{X}^{i}_{k+1}$ $=\widetilde{X}^{i}_{k}+\frac{\eta}{\sqrt{N}}\left(\delta V^{x_{i}}_{k}+\sqrt{1-\delta^{2}}\xi^{i}_{k+1}\right)-\frac{\eta^{2}}{2N}\nabla_{x_{i}}U(\theta_{k},\sqrt{N}\widetilde{X}_{k}^{i}),$
$V^{x_{i}}_{k+1}$ $=\delta^{2}V^{x_{i}}_{k}+\sqrt{1-\delta^{2}}(\delta\xi^{i}_{k+1}+\xi_{k+1}^{i,\prime})-\frac{\eta\delta}{2\sqrt{N}}(\nabla_{x_{i}}U(\theta_{k},\sqrt{N}\widetilde{X}^{i}_{k})+\nabla_{x_{i}}U(\theta_{k+1},\sqrt{N}\widetilde{X}^{i}_{k+1})).$
Finally, using $X_{k}^{i}=\sqrt{N}\widetilde{X}_{k}^{i}$, we recover,
$X^{i}_{k+1}$ $=X^{i}_{k}+\eta\left(\delta V^{x_{i}}_{k}+\sqrt{1-\delta^{2}}\xi_{k+1}\right)-\frac{\eta^{2}}{2}\nabla_{x}U(\theta_{k},X_{k}^{i})$
$V^{x_{i}}_{k+1}$ $=\delta^{2}V^{x_{i}}_{k}+\sqrt{1-\delta^{2}}(\delta\xi_{k+1}+\xi^{\prime}_{k+1})-\frac{\eta\delta}{2}\bigg(\nabla_{x}U(\theta_{k},X_{k}^{i})+\nabla_{x}U(\theta_{k+1},X_{k+1}^{i})\bigg).$
This shows that the OBABO discretisation of the rescaled diffusion, with step-size $\tilde{\eta}=\eta/\sqrt{N}$, is precisely the KIPLMC2 scheme. $\Box$
Remark A.1.
The results in Propositions A.1A.3 significantly streamline our proofs, since they allow us to connect KIPLD and our methods KIPLMC1 and KIPLMC2 to the standard underdamped diffusion and its discretisations, for which there is a rich literature of results that we can apply.
Appendix C Proofs of main results
C.1 Proof of Proposition 1
Given Proposition A.1, the system (12) has a stationary measure given by the density
$\tilde{\pi}(z,\bar{v})\propto\exp\left(-N\bar{U}_{N}(z)-\frac{1}{2}\|\bar{v}\|^{2}\right),$
@@ C.1 Proof of Proposition 1
where $\bar{U}_{N}$ is given in (11). Let us now look at $\theta$-marginal of this density, which can be written as
$\pi_{\Theta}(\theta)\propto\int e^{-\sum_{i=1}^{N}U(\theta,\sqrt{N}z_{i})-\frac{1}{2}\|\bar{v}^{z}\|^{2}}{\mathrm{d}}z_{x}{\mathrm{d}}\bar{v}.$
@@
$\pi_{\Theta}(\theta)$ $\propto\int_{\mathbb{R}^{Nd_{x}}}e^{-\sum_{i=1}^{N}U(\theta,x^{\prime}_{i})}{\mathrm{d}}x^{\prime}\propto\left(\int_{{\mathbb{R}}^{d_{x}}}e^{-U(\theta,x^{\prime})}{\mathrm{d}}x^{\prime}\right)^{N}=\exp(N\log p_{\theta}(y)),$
since $p_{\theta}(y)=\int e^{-U(\theta,x)}{\mathrm{d}}x$ by definition, where $U(\theta,x)=-\log p_{\theta}(x,y)$.
C.2 Proof of Proposition 2
Note that, by H1, $U(\theta,x)$ is jointly $\mu$-strongly convex. Let $\kappa(\theta)=-\log p_{\theta}(y)$. Using the Prékopa-Leindler inequality for strongly log-concave distributions [46, Theorem 3.8], we can see that
$\langle\theta-\theta^{\prime},\nabla\kappa(\theta)-\nabla\kappa(\theta^{\prime})\rangle\geq\mu\|\theta-\theta^{\prime}\|^{2}.$
The bound now follows using Lemma A.8 from Altschuler and Chewi [5].
C.3 Proof of Proposition 3
By Proposition A.1, the rescaled process $(\widetilde{{\mathbf{Z}}}_{t},\widetilde{{\mathbf{V}}}_{t}^{z})$ solves the standard underdamped SDE
@@ C.3 Proof of Proposition 3
${\mathrm{d}}\widetilde{{\mathbf{V}}}_{t}^{z}$ $=-\widetilde{\gamma}\widetilde{{\mathbf{V}}}_{t}^{z}{\mathrm{d}}t-\nabla N\bar{U}_{N}(\widetilde{{\mathbf{Z}}}_{t}){\mathrm{d}}t+\sqrt{2\widetilde{\gamma}}\,{\mathrm{d}}\widetilde{{\mathbf{B}}}_{t},$
with $\widetilde{\gamma}=\sqrt{N}\gamma$. By Lemmas A.2 and A.3, $\bar{U}_{N}$ is $\mu$-strongly convex and $L$-gradient-Lipschitz, hence $N\bar{U}_{N}$ is $N\mu$-strongly convex and $NL$-gradient-Lipschitz. Define
$\widetilde{\mu}:=N\mu,\hskip 20.00003pt\widetilde{L}:=NL.$
The condition $\gamma\geq\sqrt{\mu+L}$ implies
$\widetilde{\gamma}=\sqrt{N}\gamma\geq\sqrt{N(\mu+L)}=\sqrt{\widetilde{\mu}+\widetilde{L}}$.
Therefore Theorem 1 in Dalalyan and Riou-Durand [19] applies to (15) and yields, for $s\geq 0$,
@@
$W_{2}\!\left(\mathcal{L}(\widetilde{{\mathbf{Z}}}_{s}),\widetilde{\pi}_{{\mathbf{Z}}}\right)$ $\leq\;\sqrt{2}\exp\!\left(-\frac{\widetilde{\mu}}{\widetilde{\gamma}}s\right)W_{2}\!\left(\mathcal{L}(\widetilde{{\mathbf{Z}}}_{0}),\widetilde{\pi}_{{\mathbf{Z}}}\right).$
where $\widetilde{\pi}_{{\mathbf{Z}}}$ is the $\widetilde{{\mathbf{Z}}}-$marginal of $\widetilde{\pi}$ (Lemma A.1).
Now relate rescaled and original time: by definition,
$\widetilde{{\mathbf{Z}}}_{t/\sqrt{N}}=\big(\bm{\theta}_{t},N^{-1/2}{\mathbf{X}}_{t}^{1},\ldots,N^{-1/2}{\mathbf{X}}_{t}^{N}\big)^{\intercal}.$
Hence the $\theta$-component of $\widetilde{{\mathbf{Z}}}_{t/\sqrt{N}}$ is $\bm{\theta}_{t}$, and the $\theta$-marginal of $\widetilde{\pi}_{{\mathbf{Z}}}$ is $\pi_{\Theta}$. We note
$W_{2}\!\left(\mathcal{L}(\bm{\theta}_{t}),\pi_{\Theta}\right)\leq W_{2}\!\left(\mathcal{L}(\widetilde{{\mathbf{Z}}}_{t/\sqrt{N}}),\widetilde{\pi}_{{\mathbf{Z}}}\right).$
Using $s=t/\sqrt{N}$ gives
$\frac{\widetilde{\mu}}{\widetilde{\gamma}}\,\frac{t}{\sqrt{N}}=\frac{N\mu}{\sqrt{N}\gamma}\,\frac{t}{\sqrt{N}}=\frac{\mu}{\gamma}t.$
Therefore
@@
$W_{2}\!\left(\mathcal{L}(\bm{\theta}_{t}),\pi_{\Theta}\right)\leq\sqrt{2}\exp\!\left(-\frac{\mu}{\gamma}t\right)W_{2}\!\left(\mathcal{L}(\widetilde{{\mathbf{Z}}}_{0}),\widetilde{\pi}_{{\mathbf{Z}}}\right).$
Finally, with $\bar{Z}_{\star}\sim\widetilde{\pi}_{{\mathbf{Z}}}$,
$W_{2}\!\left(\mathcal{L}(\widetilde{{\mathbf{Z}}}_{0}),\widetilde{\pi}_{{\mathbf{Z}}}\right)\leq\mathbb{E}\!\left[\|\widetilde{{\mathbf{Z}}}_{0}-\bar{Z}_{\star}\|^{2}\right]^{1/2},$
which is exactly the stated bound (up to the same notation $\widetilde{{\mathbf{Z}}}_{0}\equiv{\mathbf{Z}}_{0}$ used in the proposition).
C.4 Proof of Theorem 1
By Proposition A.1, we can rewrite KIPLD as the standard underdamped diffusion with stationary measure
$\widetilde{\pi}(z,v)\propto e^{-N\bar{U}_{N}(z)-\|v\|^{2}/2}.$
We will now look at the standard kinetic Langevin Monte Carlo discretisation of the SDE and show that this scheme coincides with KIPLMC1 to utilize the bounds in Dalalyan and Riou-Durand [19] directly.
Let us write the Exponential Integrator discretisation of this scheme as we do in B.1,
$\widetilde{Z}_{n+1}$ $=\widetilde{Z}_{n}+\widetilde{\psi}_{1}^{\tilde{\eta}}\widetilde{V}_{n}^{z}-\widetilde{\psi}_{2}^{\tilde{\eta}}N\nabla\bar{U}_{N}(\widetilde{Z}_{n})+\sqrt{2\tilde{\gamma}}\zeta_{n+1}^{\prime},$ (16)
$\widetilde{V}_{n+1}^{z}$ $=\widetilde{\psi}_{0}^{\tilde{\eta}}\widetilde{V}_{n}^{z}-\widetilde{\psi}_{1}^{\tilde{\eta}}N\nabla\bar{U}_{N}(\widetilde{Z}_{n})+\sqrt{2\tilde{\gamma}}\zeta_{n+1}.$
Since the function $z\mapsto N\bar{U}_{N}(z)$ is $N\mu$ strongly convex and $NL$-gradient-Lipschitz, for $\tilde{\gamma}\geq\sqrt{N\mu+NL}$ and $\tilde{\eta}\leq\mu/(4\tilde{\gamma}L)$, Dalalyan and Riou-Durand [19, Theorem 2] directly implies that
[…]
@@ C.4 Proof of Theorem 1
where $\nu_{n}$ denotes the law of the discretisation (16) at time $n$, $\widetilde{\pi}$ the stationary measure given in Lemma A.1 and $\bar{Z}_{\star}\sim\tilde{\pi}$. We now recall from B.1, that $\tilde{\eta}=\eta/\sqrt{N}$ and $\tilde{\gamma}=\sqrt{N}\gamma$, from which we obtain,
$W_{2}(\nu_{n},\tilde{\pi})\leq C_{1}\left(1-\frac{0.75\mu\eta}{\gamma}\right)^{n}+\eta C_{2},$
where,
$C_{1}$ $=\sqrt{2}{\mathbb{E}}[\|Z_{0}-\bar{Z}_{\star}\|^{2}]^{1/2},\hskip 18.49988pt\bar{Z}_{\star}\sim\widetilde{\pi}$
$C_{2}$ $=\sqrt{2}\frac{L}{\mu}\sqrt{\frac{d_{z}}{N}}.$
Let us now recall that $\mathcal{L}(\theta_{n})$ and $\pi_{\Theta}$ are restrictions of the measures $\nu_{n}$ and $\widetilde{\pi}$ respectively. Applying this and a triangle inequality, we obtain,
$\leq W_{2}(\mathcal{L}(\theta_{n}),\pi_{\Theta})+W_{2}(\pi_{\Theta},\delta_{\bar{\theta}_{\star}})$
$\leq C_{1}\left(1-\frac{0.75\mu\eta}{\gamma}\right)^{n}+\eta C_{2}+\sqrt{\frac{C_{3}}{N}},$
where the last term is bounded by Proposition 2 with $C_{3}=d_{\theta}/\mu$.
C.5 Proof of Theorem 2
Again, we consider our rescaled standard underdamped diffusion, given by (12) and seek to directly apply the results from Monmarché [41] for our scheme.
We now write the splitting scheme as in B.2,
$\widetilde{Z}_{n+1}$ $=\widetilde{Z}_{n}+\tilde{\eta}(\tilde{\delta}\widetilde{V}^{z}_{n}+\sqrt{1-\tilde{\delta}^{2}}\xi_{n+1})-\frac{\tilde{\eta}^{2}}{2}N\nabla\bar{U}_{N}(\widetilde{Z}_{n})$ (17)
$\widetilde{V}^{z}_{n+1}$ $=\tilde{\delta}^{2}\widetilde{V}_{n}^{z}-\frac{\tilde{\eta}\tilde{\delta}}{2}(N\nabla\bar{U}_{N}(\widetilde{Z}_{n})+N\nabla\bar{U}_{N}(\widetilde{Z}_{n+1}))+\sqrt{1-\tilde{\delta}^{2}}(\tilde{\delta}\xi_{n+1}+\xi_{n+1}^{\prime}),$
where $\tilde{\delta}=e^{-\tilde{\gamma}\tilde{\eta}/2}$. Using the $\tilde{L}-$gradient-Lipschitzness and $\tilde{\mu}$ strong convexity of the function $z\mapsto N\bar{U}_{N}(z)$, we are able to apply Theorem 1 from Monmarché [41]. This result states that, for $\tilde{\gamma}\geq 2\sqrt{\tilde{L}}$ and $\tilde{\eta}\leq N\mu/(33\tilde{\gamma}^{3})$,
[…]
where $\nu_{n}$ is now the law of the discretisation (17) at time $n$, $\widetilde{\pi}$ is the stationary measure from Lemma A.1 and,
[…]
Figure 5: Multi-well Experiment. The trace plots for KIPLMC1, KIPLMC2, MPGD, IPLA and PGD on the multi-well experiment, together with density plots. We have shown in black, dashed lines the “true” values of $\bar{\theta}_{\star}$ in the trace plots and the “true” posterior density for $\theta$ in the density plots. $N=10$ and the friction and step-size are optimised for each algorithm.
Let us now recall from Theorem 2, the defined constants,
$\tilde{C}_{1}=\sqrt{3}(\sqrt{L}\lor\sqrt{L}^{-1})\mathbb{E}[\|Z_{0}-\bar{Z}^{\star}\|^{2}$ $]^{1/2},\quad\tilde{C}_{2}=\sqrt{2d_{z}}\frac{6\gamma K}{\mu}\sqrt{3}(\sqrt{L}\lor\sqrt{L}^{-1}),$
$K=L\bigg(1+e^{L\eta^{2}}$ $\left(\frac{\eta}{6}+\frac{\eta^{2}L}{24}\right)\bigg)\bigg(1+\frac{\eta L}{2\sqrt{\mu}}\bigg).$
Further, recall that, $\tilde{\mu}=N\mu$, $\tilde{L}=NL$, $\tilde{\eta}=\eta/\sqrt{N}$ and $\tilde{\gamma}=\sqrt{N}\gamma$. With this we obtain,
[…]
Again, applying the triangle inequality and Proposition 2, gives us,